Find the slope-intercept form for the line satisfying the following conditions
x-intercept 3, y-intercept 2/3

Answer by lwsshak3(11628)

Find the slope-intercept form for the line satisfying the following conditions
x-intercept 3, y-intercept 2/3
***
Equation of a line: y=mx+b, m=slope, b=y-intercept
2 points on the line: (3,0) and (0,2/3)
slope=∆y/∆x=(2/3-0)/(0-3)=(2/3)/-3=-2/9
equation: y=-2x/9+2/3